OpenStudy (anonymous):

Given f(x) = 6x^2 - x - 12 and g(x) = 2x - 3, find the function (f/g)(x).

OpenStudy (anonymous):

\[f(x)=6x^2-x-12=(2x-3)(3x+4)\] \[(f/g)(x)=((2x-3)(3x+4))/(2x-3)\] \[(f/g)(x)=3x+4\] Hope that helps :D